Calculate the volume the gas will occupy if the temperature is increased to 175


C
∘C
while the pressure is held constant.
1
Expert's answer
2020-11-17T10:41:43-0500

In this question we have 2 parameters missing which is the initial volume and initial temperature of the gas. However, let's assume a value for them. Assume that the initial volume = 200m3 and the initial temperature is 150°C. We will use Charle's law in finding the new temperature

V1/T1 = V2 /T2. . T1= 150+273 = 423K. T2 = 175 +273 = 448K

V2 = V1 ×T2 / T1 = 200×448/423 = 211.82m3